Given: Your classmate puts away $1000, $1200, $1400, $1600, three, four, five and six years, respectively after graduating from highschool. You wish to have the same amount of worth with three equal-amount savings for the zeroth, first and second year after highschool. Find: The annual amount that you have to put away.
Answer #1

Total amount saved by my classmate

= 1000 + 1200 + 1400 + 1600

= $5200

Hence if i had to collect the same amount in three years than i will have to save = 5200 ÷ 3 = $1733.33